A builder of tract homes reduced the price of a model by 15%. If the new price is $125,000, what was its original price? How much can be saved by purchasing the model?: A builder of tract homes reduced the price of a model by 15%. If the new price is $125,000, what was its original price? How much can be saved by purchasing the model?
Answer by jojo14344(880)

Remember:
OldPrice=OP
New Price=NP
NP=OP-15%OP
NP=OP(1-0.15) -------------> NP=OP(0.85)
cross(125000)147058.82/cross(0.85)=OP*cross(0.85)/cross(0.85)
$147058.82=Old Price
Savings:
=(OP-NP)/OP
=(147058.82-125000)/(147058.82)*100
=15% -----------> you're savings (as stated anyway in the problem)
Thank you,
Jojo